# A general framework for the generation of probabilistic socioeconomic scenarios: Quantification of national food security risk with application to the cases of Egypt and Ethiopia

### 07 Dec 2023 v1

@@ 3.2 Modelling food system capacity
Minimum food requirements is only the one side of the food security risk discussion, with the food system capacity (as a proxy for food supply) being its indispensable counterpart. Potential food security issues may arise when the food system capacity (in terms of calories available for human consumption) cannot cover the minimum calorie requirements of population, $C^{R}_{c,t}$, for a particular country at a certain point in time. Clearly, socioeconomic conditions, natural resources availability and of course climate conditions (directly or indirectly) affect the capacities of the national food systems. Other aspects of food security, such as physical, social and economic access to food, sustainability and nutrition, although extremely important Burlingame (2014); Garnett (2013), are bound to be subsequent considerations and at the same time policy arenas, where food security issues arising from deficient food system capacity can be mitigated or reversed. Our approach for estimating food security risk based on long-term drivers is complementary to previous attempts, that are usually focused on short-term drivers of food security (see e.g. WHO et al. (2021)). Following Ericksen’s food system framework Ericksen (2008), the main long-term drivers of food system outcomes (here calories available for human consumption) are grouped into global environmental change drivers (e.g. changes in land cover, climate variability, water and nutrient availability, etc.) and socioeconomic drivers (e.g. changes in demographics, economics, technology, etc.). The High Level Panel of Experts on Food Security and Nutrition further classifies these drivers into six categories: biophysical and environmental; technology and innovation; economic and market; political and institutional; socio-cultural; and demographic Fanzo et al. (2017).
Based on the above, to examine the effect of SSP-RCP scenarios on food security, we opt for a minimal, but reasonable and identifiable, model with a generic structure, where food system capacity is subject to the effects of the aforementioned drivers. In particular, we consider changes in socioeconomic (i.e. population, income, labour) and environmental (i.e. land, water resources, precipitation, temperature) factors that affect the food system capacity mechanism in the long-run, either directly or indirectly, through the food system activities (production, consumption, markets, trade, etc.). The above drivers are the ones for which data on future projections, consistent with each SSP-RCP scenario, are readily available or can be derived. Of course, other relevant economic drivers such as land use, labour force occupied in the food production sector, etc. could also be included Gaitán-Cremaschi et al. (2019); Fanzo et al. (2017); Van Berkum et al. (2018), but such detailed data are not usually available, especially for developing (non-OECD) countries. Also, even if available, predicting future values of such quantities under the various scenarios, and in a form compatible with the probabilistic scenarios framework, may be difficult or computationally impossible, at least with the current state of the art of economic modelling and computational resources, respectively.
To better incorporate the aforementioned features, the food system capacity is modelled by two terms corresponding to distinct contributions: (a) the domestic food production and (b) the food demand effect, which is introduced indirectly to our model through the imported and exported quantity of food. Domestic food production quantity corresponds to the total quantity of food produced within country $c$, which crucially depends on the natural resources of the country, such as land and water resources, as well as the labour force occupied in the food production sector (agriculture or livestock) and the states of the economy. On the other hand, the aspects of food demand represented by food quantity imports and exports, rely mostly on socioeconomic factors like GDP or the population level. In this perspective, the food system capacity is estimated as the aggregation of these two major contributions where other relevant quantities are determined on a bottom-up, two layer modelling approach. Historical data for food system capacity at country level are obtained by the FAO database [^18] FAO STAT Web Database: https://www.fao.org/faostat/en/#data/SCL (referred to as total food supply in the database).
In this spirit, an identifiable modelling approach for the food system capacity is represented by the two-layer scheme:
@@
  • $W_{c}$: the level of water stress of country $c$ at time $t$, i.e. the freshwater withdrawal in percentage of the available freshwater resources (according to the SDG Indicator 6.4.2191919https://www.fao.org/publications/card/en/c/CA8358EN/),
@@
The splitting of the models in two layers concerns the better presentation of the total modelling approach for food system capacity and does not introduce extra assumptions and restrictions concerning the estimation procedure. The lower layer quantities depend mainly on socio-economic drivers and potential projections for these quantities are more dependent on the various SSP scenarios. On the other hand, the upper layer quantities depend both on general socioeconomic and climate drivers, therefore potential projections on these quantities depend both on SSP and RCP scenarios. More specifically, all quantities modelled in the lower layer depend directly on population ($P$), labour ($L$) and GDP ($I$), whose evolution differs depending on the SSP scenario that is considered for the future (SSP-dependent quantities). On the other hand, the quantities modelled in the upper layer depend directly on both population ($P$) and GDP ($I$), but also on environment related quantities, such as temperature ($T$), precipitation ($Pr$) etc. whose future evolution depends on the RCP scenarios. Moreover, no interdependencies are considered between the various factors in the lower layer, while on the upper layer interdepencies between the modelled quantities are considered (for a graphical illustration of the modelling approach please see Figure 3). Extra assumptions (e.g. specific forms for evolution on agricultural labour force taking into account technological indicators) or constraints (e.g. policy or physical restrictions on using land and water resources) on the relations between the modelled quantities may be introduced by appropriate choices for the functions $f_{1},f_{2},f_{3}$ and $g_{1},g_{2},...,g_{4}$. Note that although for the case studies considered in this paper (please see Section 4) only Cobb-Douglas type models are used, these models could be substituted with any other econometric modelling approach (e.g. Translog models, CES production functions, etc) depending on the data availability.
@@
The modelling scheme (10)-(14) attempts to describe the mechanism resulting to the observed food system capacity relying on a set of socioeconomic quantities and environmental drivers. Population affects the system through various routes; it clearly affects the labour force as well as the income. Natural resources as well as environmental resources affect the domestic food production quantity. The generated food capacity for the system, allows for both domestic and imported food supply, and this depends on the level of the food quantity generated by the domestic mechanism, the effect of food demand which is captured by the food imports and exports terms, while both depend on economic quantities - GDP - which may play an important role on the ability of the country to acquire imported goods. The modelling task of these terms is performed in a regional scale, not taking into account global effects (e.g. interdependencies and interactions of the world market) which could be a direction for future work. Note also that GDP is included as a single-step delay to capture its effect on food system activities (e.g. domestic production, and related food imports and exports) but at the same time avoid reverse causal effects of these activities on GDP (Zestos and Tao (2002)). Our choice of model implicitly assumes that the drivers not included in it remain constant, while food system activities (e.g demand patterns, production technology, crop usage, food trade balance, etc.) are not disrupted (in the sense of a structural break) by the included drivers (changes in population, GDP and environmental variables, according to the SSP-RCP scenarios). As explained in the following section, under these assumptions, we are able to make predictions concerning food capacity, which will be necessary to establish a risk measure that can provide information to policy-makers, in advance for changes that need to be made in order to reduce food security risks.
@@ 3.3 A food security risk index within and across socio-economic scenarios
Based on the modelling approach presented in Sections 3.1 and 3.2, for the minimum caloric requirements ($C^{R}$) and the food system capacity ($Q^{FSC}$), we introduce in this section a new food security risk index. Combining these models we attempt to assess the food security risk at national level adopting an approach that allows for estimating/evaluating the future food security risk under different SSP-RCP scenarios, while requiring the input of less detailed data, as compared to other well established indices (such as the SDG 2.1,1 Indicator or the GFSI). The proposed indicator treats robustly the uncertainty inherent in the estimation of the key factors which affect the food security phenomenon (i.e. future population estimates which are provided in terms of probabilistic projections) either within or across the scenarios that are considered.
@@
We take this opportunity here to clarify what we mean by within and across scenarios. For various reasons (e.g. the long term horizons involved in the modelling of the phenomenon, or lack of sufficient data etc) it is not always possible for the decision maker to be aware of the exact scenario that will be materialised. Moreover, due to the long time horizon and the dynamic nature of the phenomenon, it may be that we start within the range of one scenario and in the course of time, as the phenomenon evolves, we enter within the range of a different scenario. This means that we should be able to make predictions concerning food security not only within a particular scenario (an SSP or an SSP-RCP scenario, i.e. referring to the “within scenario” case) but also taking in to account the possibility of possible multiple scenarios, or even transitions from a scenario to another. We will refer to the latter situation using the terminology “across scenarios”, i.e. the situation where SSP or SSP-RCP mixed scenarios are considered, where each scenario in the set is considered as probable with a specified probability. The rationale behind this approach, is to provide a risk assessment tool capable of supporting the decision making process of policy makers, developing a working framework that allows to take into account the effects of important drivers and the uncertainty propagated to the food security risk assessment task.
Keeping in mind the aforementioned considerations we define the new food security risk index.
Definition 3.2 (FSRI).
@@
Given the minimum caloric requirements ($C^{R}$), the food system capacity ($Q^{FSC}$) and the water stress ($W$) of a country $c$ for a specific year $t$, the Food Security Risk Index (FSRI) is defined as
@@
The proposed index expresses the percentage ratio between the minimum caloric requirements to the food system capacity capabilities at national (or regional) level, weighted by the level of water stress with respect to the sensitivity parameter $\gamma$. In the special case where the water stress risk is not taken into account (i.e. $\gamma\to 0$, assuming extremely low cost in acquiring extra water resources besides the available ones from the country, or a case of a country that will never face water stress issues), high capacity of the food system $Q^{FSC}$ comparing to $C^{R}$ leads to lower indicator values, while low $Q^{FSC}$ values comparing to $C^{R}$ should lead to increased food security risk. However, water stress risk in general significantly affects food security since most food production activities relies on water resources management and is of crucial importance, especially in areas where water resources are very limited, e.g. North Africa region. Clearly, the determination of the relevant sensitivity parameter $\gamma$ should be done with special care by the policy maker, taking into account the current and the future status of the under study area concerning the pressure on water resources and their necessity to the national or regional food production activities.
Involving water stress explicitly in the index, although is already accounted for implicitly in the other quantities $C^{R}$ and $Q^{FSC}$, contributes in its robustification with respect to potential model misspecification issues that may arise. Unavoidably, the modelling tasks of water stress and food system capacity are subject to model misspecification errors that may occur due to sudden changes in the socio-economic conditions in an area (e.g. political arrests, wars), pandemics (e.g. COVID), climate change effects (e.g. natural hazards), etc that may lead to temporally misleading estimates (overestimation or underestimation of various drivers’ effects) concerning the food demand and food supply patterns. Subsequently, poor fitting results due to such reasons, may provide unrealistic projections, i.e. enormous increase in food production quantities which may not be supported either by the capacity of the food production mechanism or the availability of natural resources (e.g. water resources). The explicit inclusion of the water stress term aims to partly treat this matter by directly introducing the risks related to the water resources availability, especially in the case where the country under study is already water stressed202020any country with water-stress indicator value above 40% is considered as highly water-stressed. According to World Resources Institute (WRI) [^21] https://www.wri.org/insights/highest-water-stressed-countries several levels for water-stress are defined with respect to the interval that the water stress indicator lies (i.e. the term $W$ in our index). In particular, the following characterizations are introduced: extremely high water stress for values greater than $80\%$, high water stress in the interval $40\%-80\%$, medium-high water stress in the interval $20\%-40\%$, low-medium water stress in the interval $10\%-20\%$ and low water stress for values less than $10\%$. The countries under study in this work, Egypt and Ethiopia, display two different patterns concerning pressure on water resources, at least for the period 2001-2018 (Figure 4). Egypt is considered as extremely high water stressed country according to the WRI definitions ($W>80\%$), and about the last two decades water stress oscillates around $120-125\%$. On the other hand, Ethiopia at early 2000s was considered a low-medium water stressed country, however it presents a clearly increasing trend and at the year 2018, the water stress is about $35\%$ characterizing the country as a medium-high water stressed one. Therefore, it seems that the pressure on water resources could provide quite different pictures even in neighbouring countries, and the inclusion of this information to the food security risk index might help in representing more accurately the actual situation.
@@
As our proposed index treats the risks in a more aggregate fashion than other detailed indices (such as SDG 2.1.1) careful and meaningful choice of the parameter $\gamma$ is required if we expect our simple index to reproduce features of other more complex indices. We must emphasise here that the simple index proposed here is suitable for projections into the future, unlike the other more complex indices. Hence, tuning the sensitivity parameter $\gamma$ in such a way such that the proposed index reproduces for past data features showing in the more complex indices, we enhance our faith in the predictive ability of our index. Given the WRI specifications related to the water stress intensity, we may provide different perspectives concerning the sensitivity of the proposed food security risk index concerning the pressure on water resources. A quite rational rule would be one to consider that the sensitivity parameter value evolves linearly (or even more steepier) with respect to the distance of the previously recorded water stress value from a safety threshold. So, if previously recorded water stress is below the safety threshold the water stress component is deactivated in the index ($\gamma=0$), otherwise $\gamma$ represents the distance from the safety threshold. For illustration purposes, we consider three different perspectives: (a) the very conservative perspective (VC) that penalizes when water stress surpasses the low-medium level ($>10\%$), (b) the less conservative perspective (LC) when the medium level ($>20\%$) is violated and (c) the no conservative perspective (NC), where only pressure beyond the high water stress threshold ($>40\%$) is taken into account and activates the relevant water risk term in the index. These perspectives allows for the determination of the sensitivity parameter in a dynamic fashion, relying on the historical water stress evolution. The perspectives introduced above are represented by the following rules:
@@
The obtained values for $\gamma$ under these three perspectives are illustrated in Figure 5 for Egypt and Ethiopia for the time period 2001-2018. Since Egypt is a extremely high water stressed country, all three rules lead to strictly positive values for $\gamma$ of that evolve on exactly the same manner with different horizontal position (translated lower or higher depending on the perspective). On the other hand, for Ethiopia the NC perspective does not allocate any positive value on $\gamma$ since the water stress index for this period is below the activation threshold of $40\%$. The LC perspective provides positive values for $\gamma$ after the year 2004 with increasing trend, while under the VC perspective positive values are allocated on $\gamma$ for the whole period with also increasing trend.
@@
Although even more sophisticated approaches can be considered in the method that the two risks are aggregated, it seems that this quite simple threshold-based approach, allows to qualitatively approximate well established indices like PoU (SDG 2.1.1) that require more detailed data. This fact is illustrated in Figure 6 where the PoU indicator and the proposed index (FSRI) for different values of $\gamma$ are depicted for the both countries under study. Note that for comparison purposes, the case where water stress is not taken into account to the food security risk calculation ($\gamma=0$) is also included. Clearly, for Egypt the tendency of the PoU indicator is better represented by FSRI for any of the choices that allocate positive values on $\gamma$ comparing to the case where the term of water risk is excluded from the calculation ($\gamma=0$). Moreover, for Ethiopia the nice behaviour of the FSRI in approximating the tendency of PoU is also observed without any major misspecifications for any choice of $\gamma$. However, this is expected to change if the country enters the highly water-stressed stage, where the water risk component of the index is expected to play more important role in the better representation of the food security risk situation.

### 14 Dec 2023 v2

@@ 3.2 Modelling food system capacity
Food security risk is influenced not only by minimum food requirements but also critically by the capacity of national food systems. A potential risk emerges when a country’s food system capacity, defined in terms of available calories for human consumption, fails to meet its population’s minimum caloric requirements $C^{R}_{c,t}$. The Food and Agriculture Organization (FAO) succinctly defines the concept of food self-sufficiency [^14] FAO (1999) https://www.fao.org/3/X3936E/X3936E00.htm as ”the extent to which a country can satisfy its food needs from its own domestic production”. However, self-sufficiency is not a synonym of food security in the sense that a nation’s strategy may not solely rely on domestic production. Global markets always function and play a crucial role in ensuring food availability, especially in scenarios where local production is insufficient or disrupted. Thus, the capacity of a food system is significantly influenced not only by domestic factors but also by the country’s economic engagement with global markets. This includes the planning of economic sectors involved in food production and distribution, choice of crops, and crucially, the dynamics of imports and exports. Acknowledging this broader perspective is essential for a comprehensive understanding of national food security and its reliance on both domestic supply and international market interactions. Socioeconomic conditions, natural resources, and climate factors all play pivotal roles in shaping the capacity of these food systems. While other dimensions of food security, such as physical, social, and economic access to food, sustainability, and nutrition, are undeniably crucial Burlingame (2014); Garnett (2013), they are often subsequent considerations and at the same time policy areas for addressing issues stemming from inadequate food system capacity.
Our method for assessing food security risk is designed to complement existing models, which often focus on immediate, short-term factors WHO et al. (2021). In line with Ericksen’s food system framework Ericksen (2008), we identify two primary groups of long-term drivers influencing food system outcomes: global environmental change drivers, including variables such as land cover changes, climate variability, and water and nutrient availability; and socioeconomic drivers, encompassing factors like demographics, economic conditions, and technological progress. Additionally, the High Level Panel of Experts on Food Security and Nutrition expands on this classification, categorizing these drivers into six broader areas: biophysical and environmental; technology and innovation; economic and market; political and institutional; socio-cultural; and demographic Fanzo et al. (2017). This comprehensive approach enables a more holistic understanding of the multifaceted factors impacting food security over extended periods.
To evaluate the implications of SSP-RCP scenarios on food security, we utilize a broadly structured model that reflects the influences of key drivers on the capacity of the food system. Our focus is on how shifts in socioeconomic elements (like population dynamics, income levels, and labor force) and environmental factors (such as land and water resources, along with changes in precipitation and temperature) shape the food system over time. These variables play critical roles in food-related activities, including production, consumption, market dynamics, and trade. The model aligns with available or projectable data for these factors under each SSP-RCP scenario. However, the inclusion of additional economic drivers, for instance, land use patterns and the workforce engaged in food production, while valuable Gaitán-Cremaschi et al. (2019); Fanzo et al. (2017); Van Berkum et al. (2018), faces practical limitations. Detailed data on these aspects are often not accessible, especially for non-OECD countries. Furthermore, forecasting these variables for future scenarios poses significant challenges, both in terms of the complexity of economic modeling and the computational resources required. To effectively model the food system capacity, our approach integrates a balanced interplay of domestic supply, imports, exports, and inferred demand. The first component, domestic food production, is contingent upon a country’s natural resources like land and water, the labor force in agriculture or livestock sectors, and the prevailing economic conditions. The second component, while not directly measuring food demand due to data limitations, uses domestic food supply, imports, and exports as proxies to approximate demand, underpinning the equilibrium state of the food system’s capacity. The latter part of the model acknowledges that food system capacity observed is an equilibrium between the amount of food a country produces, the food it imports and exports, and the overall demand within the country. Historical data on food system capacity are sourced from the FAO database [^15] FAO STAT Web Database: https://www.fao.org/faostat/en/#data/SCL (referred to as total food supply in the database).
In 10-14, we represent that model in a two-layer structure. This structure is primarily a representational tool, designed to enhance clarity and organization in the presentation of the model. It is important to clarify that these layers do not imply any hierarchical or sequential order in the estimation process, nor do they introduce any additional complexity or assumptions into the modeling approach. In essence, the division into ’lower’ and ’upper’ layers is a conceptual framework used for better illustrating how different types of drivers—socioeconomic and climate-related—affect the food system capacity. The lower layer, focuses on socioeconomic drivers, and the upper layer, encompassing both socioeconomic and environmental factors. This distinction is made to aid in the understanding and explanation of the model, rather than to suggest a layered approach in the actual computation or estimation of food system capacity. All factors, regardless of the layer they are categorized in, are integral to the model and contribute collectively to the estimation of food system capacity and, by extension, to the assessment of food security risk. Under the aforementioned considerations, we set the two-layer model:
@@
  • $W_{c}$: the level of water stress of country $c$ at time $t$, i.e. the freshwater withdrawal in percentage of the available freshwater resources (according to the SDG Indicator 6.4.2161616https://www.fao.org/publications/card/en/c/CA8358EN/),
@@
In our model, the quantities in the lower layer are directly influenced by population ($P$), labor ($L$), and GDP ($I$), with their trajectories varying according to the specific SSP scenario being considered (SSP-dependent quantities). Conversely, in the upper layer, the quantities are not only affected by population ($P$) and GDP ($I$) but also by environmental variables like temperature ($T$) and precipitation ($Pr$), which are influenced by the chosen RCP scenarios. While the lower layer assumes no interdependencies among its various factors, the upper layer does account for interactions between the modeled quantities (a graphical representation of this modeling approach can be found in Figure 3). Furthermore, additional assumptions or constraints on the relationships between these quantities can be implemented through suitable choices of functions $f_{1},f_{2},f_{3}$ for the lower layer and $g_{1},g_{2},...,g_{4}$ for the upper layer. These might include, for example, specific evolutionary forms for the agricultural labor force considering technological changes, or policy or physical constraints on land and water resource usage. It is important to note that, although the case studies in this paper (refer to Section 4) employ Cobb-Douglas type models, the model is flexible enough to accommodate other econometric approaches such as Translog models or CES production functions, depending on data availability.
@@
The modeling framework outlined in (10)-(14) seeks to delineate the processes leading to the observed capacity of the food system, drawing on a range of socioeconomic and environmental drivers. Population plays a multifaceted role, influencing both labor force and income. Similarly, natural and environmental resources are key determinants of domestic food production. The food system’s capacity emerges from the equilibrium between domestic food production and the interplay of food imports and exports. This balance reflects a country’s participation in the global food market and its economic capability, influenced by GDP, to augment domestic production and satisfy demand. In our approach, these factors are modeled at a regional scale, without incorporating global influences such as market interdependencies, which presents an avenue for future exploration. Additionally, GDP is introduced with a one-step delay in the model to capture its effect on food system activities like domestic production and trade, while simultaneously mitigating potential reverse causal impacts of these activities on GDP (Zestos and Tao (2002)). In the following section, we will discuss how these foundational assumptions are instrumental in enabling projections of food capacity that are vital for crafting a risk measure. This measure is designed not just to signal vulnerabilities but to also catalyze preemptive strategies that could fortify a nation’s food security stance.
Our model operates under the implicit assumption that the food system will adapt gradually to changes in key drivers, as outlined by SSP-RCP scenarios, without experiencing abrupt structural shifts in activities such as demand patterns, production technology, crop usage, and food trade balance. Furthermore, it assumes that factors not explicitly modeled will remain up to time trend, captured by $t$. This approach, while simplifying the complexity of real-world dynamics, strategically identifies variables that policymakers have the capacity to influence. As detailed in the ensuing section, this modeling choice allows policymakers to utilize the food security risk index as a diagnostic tool. Should the index reveal areas of concern, the unmodeled drivers offer a set of adjustable parameters through which policymakers can enact changes to strengthen food security measures
@@ 3.3 A food security risk index within and across socio-economic scenarios
Concerning the task of quantifying food security risk, several indicators have been introduced in the literature so far. The United Nations has introduced the SDG Indicator 2.1.1, also known as the Prevalence of Undernourishment (PoU) indicator. This indicator is defined [^17] https://unstats.un.org/sdgs/metadata/?Text=&Goal=2&Target as ‘an estimate of the proportion of the population whose habitual food consumption fails to provide the necessary dietary energy levels for maintaining a normal active and healthy life, expressed as a percentage’. However, the PoU indicator has faced criticism due to the extensive and detailed data requirements for its computation. It necessitates periodic household surveys, comprehensive information on food acquisition per household, dietary habits, and additional data on the total available food for human consumption to correct potential biases. Consequently, calculating this indicator is a complex task, requiring granular household-level information. Furthermore, forecasting future trends or immediate changes in this index poses significant challenges due to these data requirements.
Also, FAO’s Food Insecurity Experience Scale (FIES) serves as an important tool in assessing food security, offering internationally comparable estimations of the extent of food access difficulties encountered by individuals and households. [^18] https://www.fao.org/policy-support/tools-and-publications/resources-details/en/c/1236494/ It contributes to the monitoring of progress towards Sustainable Development Goal (SDG) Target 2.1, which is dedicated to ending hunger and ensuring universal access to food. The FIES derives its insights from direct interviews, quantifying the severity of food insecurity faced by people, and is instrumental in measuring the advances towards achieving SDG Target 2.1. This measure supplements the insights provided by the SDG Indicator 2.1.1, the Prevalence of Undernourishment (PoU) indicator, by capturing a more detailed picture of food access challenges.
Another prominent index is the Global Food Security Index (GFSI) [^19] https://impact.economist.com/sustainability/project/food-security-index/, which evaluates food affordability, availability, quality, safety, sustainability, and adaptation across 113 countries. The GFSI is a dynamic benchmark model, both quantitative and qualitative, comprised of 68 unique indicators that measure the drivers of food security in various countries, spanning both the developed and developing world. The complexity of estimating the GFSI surpasses that of SDG 2.1.1 due to its comprehensive nature, which aims to reflect the state of global food security. As a result, its utility lies primarily in presenting the current state of affairs rather than projecting future scenarios, given its intricate construction and wide-ranging scope.
Drawing upon the methodologies outlined in Sections 3.1 and 3.2, which detail the calculations for minimum caloric requirements ($C^{R}$) and food system capacity ($Q^{FSC}$), we present a novel food security risk index in this section. This index is the result of an integrated approach, combining the aforementioned models to assess national-level food security risks. Our methodology is adept to evaluate future food security risks under various SSP-RCP scenarios, with the added benefit of requiring less detailed data inputs compared to established indices like the SDG 2.1.1 and FIES Indicators or the GFSI. A key strength of our proposed index is its robust treatment of uncertainty, particularly in the estimation of critical factors influencing food security. This includes addressing uncertainties in future population estimates, which are provided as probabilistic projections in our model.
Building on the methodologies delineated in Sections 3.1 and 3.2, which elaborate on the computation of minimum caloric requirements ($C^{R}$) and food system capacity ($Q^{FSC}$), this section introduces a new index for assessing food security risk. This index emerges from an amalgamation of the aforementioned models, facilitating the assessment of food security risks at a national level. Our model is adept at projecting future food security risks within diverse SSP-RCP scenarios, with the distinct advantage of minimizing the need for detailed data inputs, unlike more conventional indices such as the SDG 2.1.1 and FIES Indicators or the GFSI. A prominent feature of our proposed index is its rigorous approach to managing uncertainty, especially in the estimation of pivotal factors that impact food security. This involves incorporating uncertainties in future population figures, which are rendered through probabilistic projections in our model.
@@
Our index is designed to be flexible and reliable, capable of assessing food security risks both within individual scenarios and across a range of different SSP-RCP scenarios. We take this opportunity here to clarify what we mean by within and across scenarios. For various reasons (e.g. the long-term horizons involved in the modelling of the phenomenon, or lack of sufficient data etc) it is not always possible for the decision maker to be aware of the exact scenario that will be materialised. Moreover, due to the long time horizon and the dynamic nature of the phenomenon, it may be that we start within the range of one scenario and in the course of time, as the phenomenon evolves, we enter within the range of a different scenario. This means that we should be able to make predictions concerning food security not only within a particular scenario (an SSP or an SSP-RCP scenario, i.e. referring to the “within scenario” case) but also taking in to account the possibility of possible multiple scenarios, or even transitions from a scenario to another. We will refer to the latter situation using the terminology “across scenarios”, i.e. the situation where SSP or SSP-RCP mixed scenarios are considered, where each scenario in the set is considered as probable with a specified probability. The rationale behind this approach, is to provide a risk assessment tool capable of supporting the decision making process of policy makers, developing a working framework that allows to take into account the effects of important drivers and the uncertainty propagated to the food security risk assessment task.
@@
Keeping in mind the aforementioned considerations we define the new food security risk index. Given the minimum caloric requirements ($C^{R}$), the food system capacity ($Q^{FSC}$) and the water stress ($W$) of a country $c$ for a specific year $t$, the Food Security Risk Index (FSRI) is defined as
@@
The proposed index expresses the percentage ratio between the minimum caloric requirements to the food system capacity at national (or regional) level, weighted by the level of water stress with respect to the sensitivity parameter $\gamma$. In the special case where the water stress risk is not taken into account (i.e. $\gamma\to 0$, assuming extremely low cost in acquiring extra water resources besides the available ones from the country, or a case of a country that will never face water stress issues), high capacity of the food system $Q^{FSC}$ comparing to $C^{R}$ leads to lower indicator values, while low $Q^{FSC}$ values comparing to $C^{R}$ should lead to increased food security risk. However, water stress risk in general significantly affects food security since most food production activities relies on water resources management and is of crucial importance, especially in areas where water resources are very limited, e.g. North Africa region. Clearly, the determination of the relevant sensitivity parameter $\gamma$ should be done with special care by the policy maker, taking into account the current and the future status of the under study area concerning the pressure on water resources and their necessity to the national or regional food production activities.
Incorporating water stress explicitly into our food security risk index, despite its implicit consideration within the quantities $C^{R}$ and $Q^{FSC}$, enhances the model’s robustness against potential misspecification issues. The modeling of water stress and food system capacity is inherently prone to errors due to unexpected socio-economic shifts (e.g., political unrest, wars), pandemics (like COVID-19), or climate change impacts (such as natural disasters), which can lead to transiently inaccurate estimates for food demand and supply patterns. These errors might result in unrealistic projections, such as an overestimated increase in food production that surpasses the limits of production mechanisms or natural resource availability, particularly water resources. The explicit inclusion of water stress in the model addresses this issue by directly considering the risks associated with water resource availability, especially in regions already experiencing high water stress [^20] Countries with a water-stress indicator value above 40% are considered as highly water-stressed. According to the World Resources Institute (WRI) [^21] https://www.wri.org/insights/highest-water-stressed-countries, water stress is classified into several levels: extremely high water stress for values over 80%, high water stress between 40%-80%, medium-high water stress from 20%-40%, low-medium water stress between 10%-20%, and low water stress below 10%. In our case studies, Egypt and Ethiopia, each country exhibits distinct water resource pressures for the period 2001-2018 (Figure 4). Egypt falls into the extremely high water stress category (WRI definition, $W>80\%$) with water stress levels fluctuating around 120-125% in the last two decades. Ethiopia, initially a low-medium water stressed country in the early 2000s, has shown an increasing trend, reaching approximately 35% water stress in 2018, categorizing it as medium-high water stressed. The divergent scenarios in these neighboring countries highlight the importance of including water stress data in the food security risk index to more accurately reflect the actual situation.
@@
Our proposed index aggregates risks more broadly compared to detailed indices like SDG 2.1.1, necessitating a careful and meaningful selection of the parameter $\gamma$. It is important to note that our simpler index is particularly suited for future projections, a capability not as readily available in more complex indices. By calibrating the sensitivity parameter $\gamma$ to replicate features observed in these complex indices based on past data, we enhance confidence in the predictive potential of our index. According to the World Resources Institute (WRI) specifications for water stress intensity, our index can offer varied insights into the sensitivity to water resource pressures. A logical approach would be to adjust the value of $\gamma$ in relation to the deviation of the previously recorded water stress value from a defined safety threshold. Specifically, if the past water stress value falls below this threshold, the water stress component in the index would be inactivated ($\gamma=0$); otherwise, $\gamma$ would represent the degree of deviation from the threshold.
For illustrative purposes, we adopt three distinct perspectives: (a) a very conservative (VC) stance, where $\gamma$ is adjusted when water stress exceeds the low-medium level ($>10\%$), (b) a less conservative (LC) approach, responding to breaches of the medium level ($>20\%$), and (c) a non-conservative (NC) viewpoint, which only considers significant deviations beyond the high water stress threshold ($>40\%$). These perspectives enable dynamic determination of the sensitivity parameter, reflecting the historical evolution of water stress and can be represented by the following rules:
@@
The calculated values for $\gamma$ under these three perspectives (VC, LC, NC) are depicted in Figure 5 for Egypt and Ethiopia over the period 2001-2018. For Egypt, which is classified as an extremely high water-stressed country, all three rules result in strictly positive values for $\gamma$. These values evolve in a similar manner for each perspective, differing only in their horizontal positioning, which varies according to the specific rule applied. In the case of Ethiopia, the NC perspective assigns no positive value to $\gamma$, as the water stress index remains below the $40\%$ activation threshold throughout this period. Conversely, the LC perspective begins to assign positive values to $\gamma$ after 2004, exhibiting an upward trend. Under the VC perspective, $\gamma$ is allocated positive values for the entire period, also following an increasing trend.
@@
While more intricate methods could be employed in aggregating the two risks, our straightforward, threshold-based approach seems to effectively approximate established indices like the PoU (SDG 2.1.1), which necessitate more detailed data. This is demonstrated in Figure 6, where we compare the PoU indicator with our proposed Food Security Risk Index (FSRI) using different values of $\gamma$ for both countries under study. For the sake of comparison, we also include scenarios where water stress is not factored into the food security risk calculation ($\gamma=0$). Notably, for Egypt, the trend of the PoU indicator is more accurately mirrored by the FSRI when positive values are assigned to $\gamma$, as opposed to the scenario excluding the water risk component ($\gamma=0$). In the case of Ethiopia, the FSRI closely approximates the trend of the PoU without significant discrepancies, regardless of the $\gamma$ value chosen. However, this alignment might shift if Ethiopia transitions into a highly water-stressed stage, at which point the water risk component of the index would likely become more influential in accurately depicting the food security risk.